OpenStudy (anonymous):

Calculate the volume of a sphere where the radius is 9 meters A. 81 cubic meters B. 324 cubic meters C. 972 cubic meters D. 2916 cubic meters

OpenStudy (stamp):

I gave you the equation earlier, quit being a lazy nuisance and do your own work. Here it is again, and actually do it this time please. Ask questions if you need further help, do not make another thread hoping somebody will spoon feed you the answer and wipe off your lips. \[V=\frac{4}{3}\pi r^3\] \[r=9m\] \[V=\frac{4}{3}\pi (9m)^3\]

OpenStudy (anonymous):

why dont you kiss my retrice

OpenStudy (stamp):

@matthewenit This is why you are asking for help, because you are immature and do not understand what it means to work for something.

OpenStudy (anonymous):

exactly why im doing geometry in 8th grade?

OpenStudy (stamp):

\[V=\frac{4}{3}\pi (9m)^3\]

OpenStudy (stamp):

@matthewenit Do not get arrogant, you are still a child.

OpenStudy (anonymous):

of course im still a child, thats why im immature. :)

OpenStudy (anonymous):

339 is what i got,..

OpenStudy (stamp):

\[V=\frac{4}{3}\pi (9^3)=\frac{4}{3}\pi*9*9*9\]Divide the 3 into one of the 9s\[4\pi\ 3(9)(9)=12\pi81=960+12=972\pi\ m^3\]

OpenStudy (anonymous):

oh i wasnt cubing the 9 i was squaring it. thanks sorry for being a nuisance.

OpenStudy (stamp):

Thank you.